Find the inverse of each function, if it exists. Show your work and circle your answer.
f(x)=10x cubed

Answer by CubeyThePenguin(3113)

Solve for x.

y = 10x^3
y/10 = x^3
cbrt(y/10) = x

f^(-1)(x) = cuberoot(x/10)